A closed rectangular box whose length is double its width has a total surface area of 600 cm². Find the dimensions of the box with maximum volume.


1
Expert's answer
2021-10-04T16:31:47-0400

The surface area of a rectangular prism with the width "w", length "l", and height "h" h is given by:

"S=2(wl+wh+lh)"

Substitute "S=600, ~\text{and}~~l=2w" into the above equation:

"600=2(w \cdot 2w +wh+2wh)"

Simplify the equation:

"600=2(2w^{2} +wh+2wh)"

Find "h" in terms of "w":

Divide both sides of the equation by "2":

"300=2w^{2} +wh+2wh"

Combine like terms:

"300=2w^{2} +3wh"

Move the expression "3wh" to the left-hand side and change its sign:

"300-3wh=2w^{2}"

Move the expression to the left-hand side and change its sign:

"-3wh=2w^2-300"

Divide both sides of the equation by "-3w":

"h=\frac{2w^2-300}{-3w}"

The volume of the rectangular box is given by:

"V=w \cdot l \cdot h"

Substitute "l=2w" and "h=\frac{2w^2-300}{-3w}" into the equation:

"V=w \cdot 2w \cdot \frac{2w^2-300}{-3w}"

Simplify the equation:

"V=-\frac{4}{3}w^{3}+200w"

To find the maximum volume, differentiate "V" with respect to "w"

"\frac{dv}{dt}=-4w^{2}+200"

Set "\frac{dv}{dt}=0"

"-4w^{2}+200=0"

Move the expression "200" to the left-hand side and change its sign:

"-4w^{2}" "=-200"

Divide both sides by "-4" :

"w^2=50"

Take the square root for both sides:

"w=\pm 5\sqrt{2}"

Since there is no negative lengths, exclude the negative sign:

"w=5\sqrt{2}"

"l=2w=2 \times 5\sqrt{2}=10\sqrt{2}"

"h=\frac{2(5\sqrt{2})^2-300}{-3(5\sqrt{2})}"

Using a calculator:

"h=\frac{20\sqrt{2}}{3}"